Minutes of the 17th SOHO SWT Meeting

Goddard Space Flight Center, Greenbelt, MD 7–8 December 1995 Contents

1 Action Items SOHO SWT 17 2 2 Agree Agenda and Actions Revision 2 3 Maneuvre Plans for the Remainder of the Mission 2 4 Status of Experiments 2 5 Commissioning and Early Operations 3 6 SOT Concept 4 7 PR Matters 4 8 Future Workshops and Colloquia 4 9 SOHO Bibliography 5 10 AOB 5 Annex 1: Agenda 7 Annex 2: List of Participants 8 Annex 3: Congratulations from the Yohkoh Team 11 Annex 4: Maneuver Plans for the Remainder of the Mission 12 Annex 5: Commissioning Status and Plans 18 Annex 6: Line of Communication during Commissioning 22 Annex 7: SOHO Science Report Issue 10 26 Annex 8: Layout of Press Releases 33 Annex 9: SOHO Science Related Conferences 35

1 Action Items SOHO SWT 17

Action 17-1: on PS to clarify issue of experiment movements during PTCS calibration. (already closed by time of writing of minutes: meeting scheduled for Wednesday, Dec 13) Action 17-2: on PIs to provide power-off recovery procedures to PS before Dec 15. Action 17-3: on PS to clarify with FOT execution of emergency switch-off procedures on direct PI request. (already closed by time of writing of minutes; see memo “Line of communication for PIs during commissioning” from 11 December 1995) Action 17-4: on PIs to provide e-mail addresses of SPWG members. Input to B.Fleck before Dec 15. Action 17-5: on PS to contact R.Harrison and arrange free block for SOHO presentations at EGS (Dec 15). (already closed by time of writing of minutes) Action 17-6: on PS to arrange special SOHO session at 1996 SPD meeting (Dec 15). Action 17-7: on PIs to provide bibliography information (list of papers with abstracts, database used). Input to B.Fleck before January 30. Action 17-8: on PIs to provide info on space requests (EOF + EAF) to A.Poland (Dec 15).

2 Agree Agenda and Actions Revision

Agenda — annex 1 List of Participants — annex 2 V. Domingo welcomed all participants, commended all parties involved for the most successful beginning of the mission and forwarded the congratulations of the Yohkoh team on the successful launch of the SOHO mission (annex 3).

Actions revision

Action 15.1 on PIs closed to comment on present commissioning plan.

Action 15.2 on PIs closed to provide manning plan for EOF/EAF including names. Input to PS before August 31.

Action 15.3 on PIs closed to ensure that the relevant steps (software changes etc.) are taken to comply with the adoption of UTC at the Earth as the time reference for data.

Action 15.4 on PS closed to organize meeting of US and European archive representatives to arrange schedules and details for archive development.

3 Maneuvre Plans for the Remainder of the Mission

Becky Short presented the maneuvre plans for the remainder of the mission at the daily 9 am status meeting on Friday morning, Dec 8 (see annex 4). Conclusion: HOI and MCC2 can be optimized by performing larger maneuver at MCC2 which then becomes an Orbit Shaping Maneuver (OSM). Date of MCC2 estimated: first week of January.

4 Status of Experiments

1. GOLF was switched on, all temperatures fine, now off 2. VIRGO

very smooth switch on on December 6, cover released and first measurements made

3. SOI/MDI turned on for CCD bake-out; CCD at 30^{∘}C; turned on filter and oven heater, rest will be turned on on Monday, Dec 11

4. SUMER front 6^{∘} warmer than expected, but not critical; expected to vanish during operations when instrument is controlable

5. CDS still switched off

6. LASCO/EIT still switched off

7. UVCS switched on, door cracked, launch lock released, front part warmer than expected, but not critical

8. SWAN just turning on

9. CELIAS still off; temperatures o.k.

10. COSTEP/ERNE shortly prior to switch on, temperatures normal

5 Commissioning and Early Operations

P. Martens presented a high level commissioning plan for the whole commissioning phase and the commissioning status as of December 6 (see annex 5). A. Gabriel supported R. Harrison’s request to have a detailed strawman plan for January/February. P. Martens expressed his doubts that such a plan would have any use since at this time it is fairly difficult to come up with a realistic schedule for the Jan/Feb time frame. There is an action on the Project Scientist to clarify the issue of experiment movements during PTCS calibration (Action 17-1). It was clarified that the CELIAS switch on will not necessarily monopolize the NRT channel for 10 days; other activties can be run in parallel. F. Felici pointed out, though, that he would like to keep the principle that first time actions of instruments should not be run in parallel with other activities, but only run in dedicated sessions. This has to be discussed on a case by case basis. Project and SOC will get together to discuss the cases which promise to be most beneficial to save time.

Switch on order after power-off: GOLF, VIRGO, MDI, SUMER, CDS, EIT/LASCO, UVCS, SWAN, CELIAS, CEPAC with TSTOL procedures subsequently run by the FOT for all the experiments, then shared NRT session. SUMER stressed that the substitution heaters must be switched on separately by the FOT, not collectively. There is an action on the PIs to provide power-off recovery procedures to the PS by December 15 (Action 17-2). UVCS does not see a contamination risk by thruster firing after an ESR and therefore does not want to close door in case of an ESR.

Line of communication during commissioning: V. Domingo presented the line of communication during commissioning (see annex 6). J. Kohl raised the issue of an emergency shut down. There is an action on the Project Scientist to clarify with the FOT the execution of an emergency switch-off procedure on direct PI request, without the necessity of going through the normal line of communication (Action 17-3).

6 SOT Concept

A. Gabriel explained his concerns about the present SOT concept from the viewpoint of a remote PI. After some discussion it was agreed to have monthly meetings of the SOHO Science Planning Working Group (SPWG), in addition to the quartely SWT meetings. The monthly SPWG meetings replace the monthly SOT meetings. At these meetings the three-month plans will be refined, JOPs and coordinated observations with ground-based observatories and other spacecraft discussed. A. Gabriel also suggested to establish a SOHOSPWG e-mail distribution list at Stanford. There is an action on the PIs to provide the e-mail addresses of the SPWG members of their teams. Input to B. Fleck before Dec 15 (Action 17-4).

7 PR Matters

SOHO Science Report B. Fleck presented the most recent issue of the SOHO Science Report (Issue 10; see annex 7). For the future it is planned to have the SOHO Science Reports issued quarterly, with inputs from all the experiment teams highlighting some of the most recent scientific discoveries. The PIs are requested to bring their input to the quarterly SWT meetings where the content of the Science Report will be discussed. V. Domingo will contact ESA’s Publication Division to seek their support in producing forthcoming issues of the SOHO Science Report. If possible, they should be done in color.

Layout of Press Releases The layout of future SOHO Press Releases was briefly discussed. It was agreed to generally follow the Hubble/STSI approach (see annex 8), i.e. the backside of the photographs should include the SOHO logo with ESA and NASA labels, title of the release, description of figure, and give credit to the scientist(s) and institutes providing the content.

SOHO Press Day Given the long list of major scientific meetings starting May 6 with the EGS General Assembly in The Hague, V. Domingo proposed to hold the big “First month of SOHO Science” media event on May 2, 1996. All the PIs — and the PRWG members in particular — are requested to think about “science stories” for this press day. Ideas should be submitted to the Project Scientists.

8 Future Workshops and Colloquia

B. Fleck presented a list of SOHO science related conferences and some background information on the EGS sessions (see annex 8). Several PIs expressed their uneasiness to submit abstracts before having seen first results from their instruments. There is an action on the Project Scientist to contact R.Harrison and arrange a free block for SOHO presentations at the EGS (Action 17-5).^{1} The PIs also felt that there are simply too many meetings to attend all of them. It was therefore suggested to appoint “ambassadors” to give SOHO overview presentations at the numerous spring 1996 meetings. Main events for the SOHO community to show first and early results: EGS, SPD, COSPAR, and fall AGU. There is an action on the Project Scientist to arrange a special SOHO session at 1996 SPD meeting (Action 17-6). E. Antonucci presented the various events at COSPAR 96. It was agreed to have 1 invited talk by each experiment team. E. Antonucci to send invitations. There will be a special session “First Results from the Solar and Heliospheric Observatory” at the fall 1996 AGU meeting. Convenors: V. Domingo, A. Poland, A. Gabriel and G. Brueckner. M. Huber will contact the organizers of the 8th European Meeting on Solar Physics in Thessaloniki to arrange for a 2 hour open SOHO session. V. Domingo will organize a SOHO session and/or overview talk at the General Assembly of the European Physical Society in Seville. B. Fleck will investigate boundary conditions and constraints for Symposia or Joint Discussions during the 1997 IAU General Assembly in Kyoto. It is planned to organize an ESLAB Synposium in the second half of 1997 as follow up of the 1985 ESA Workshop on Future Missions in Solar, Heliospheric & Space Plasma Physics in Garmisch-Partenkirchen (ESA SP-235).

1Session ST6.01: “Workshop on Results from the Solar Terrestrial Science Programme” is open, so people can simply come along with any data they wish to show and present.

9 SOHO Bibliography

The SOHO Project Scientist Team is maintaining a comprehensive list of SOHO related publications, published regularly in the SOHO Science Reports. The list is based on re- and preprints received, and the INSPEC and NASA database of IRS at ESRIN/ESA. Presently, this list is kept as a simple Latex text file. It is planned to establish an interactive database system at the SOHO archive which will contain a complete list of all SOHO papers, if possible including abstracts. L. Sanchez will circulate a questionnaire among the PI teams to gather information about the systems presently used. There is an action on the PIs to provide bibliography information (list of papers with abstracts, database used). Input to B.Fleck before January 30 (Action 17-7).

10 AOB

MCE Huber, the Head of ESA’s Space Science Department expressed his thanks to the PIs’ for their collaboration in the PR activities surrounding launch, and especially in the corporate TV broadcast. He had had favourable reactions from people who had attended the transmission at ESA Head Office in Paris. (Unfortunately, there were technical problems at ESTEC, which prevented the projection of the broadcast.) He also congratulated the PIs on the successful start of the misssion, and conveyed the congratulations of ESA’s Science Programme Committee, which had met earlier in the week in Paris. SPC discussed the international aspect of the SOHO mission and supports giving proper credit to ESA and NASA. The Head of Space Science Department also announced a visit by the ESA Director Science, scheduled for 8 February, 1996. M. Huber mentioned that the Director had asked him for a report on the findings on the present visit, in particular as far as the lines of authority for operations and proper support of the international aspect of the mission was concerned. The Director had the intention to formally appoint the Head of Space Science Department to coordinate the latter task, and to confirm this appointment with the Associate Administrator (W. Huntress) and the Director of GSFC (J. Rothenberg). M. Huber asked the PIs for their support in this delicate task. He enumerated a number of elements needed to achieve properly balanced credit. Some of these had in fact just been discussed in the preceding agenda item (representation of SOHO at conferences). One would probably also have to explore whether further measures (perhaps even a “mild” renaming of the mission) might be used to make the public understand the international nature of the SOHO mission.

L. Sanchez will establish and maintain a SOHO status page on the Web with input from the PIs and the Project. To improve the information flow, H. Kunow suggested to post meetings and changes thereof at the bulletin boards in the EOF and EAF. J. Kohl requested to arrange for a SIC meeting adjacent to next SWPG meeting. L. Sanchez mentioned that there is now detailed information on the SOHO archive available. The Project Scientists will contact Science, Nature, Scientific American etc. regarding special SOHO papers. There is an action on the PIs to provide info on space request (EOF + EAF) to A.Poland (Action 17-8).

Next Meeting It was agreed to have the monthly SPWG meetings on the second Friday of each month, i.e. first SPWG meeting on January 12. Next SWT meeting: 8-9 February 1996 (together with SPWG meeting).

Annex 1: Agenda

7

Annex 2: List of Participants

8

Annex 2: Congratulations from the Yohkoh Team

11

Annex 4: Maneuver Plans for the Remainder of the Mission

12

Annex 5: Commissioning Status and Plans

18

Annex 6: Line of Communication during Commissioning

22

Annex 7: SOHO Science Report Issue 10

26

Annex 8: Layout of Press Releases

33

Annex 9: SOHO Science Related Conferences

35